I need help with solving these two questions
Thanks
1. The equilibrium constant for the following reaction at 100°C is 0.11. If M ethanol (C2H,OH) and 0.810 M acetic acid (CH3COOH) are combined, what is the equilibrium concen- tration of ethyl acetate (CH3CO2C2Hs)? 6.816 2H,OH (aq) +CH,COOH(ag)CH,CO,C2Hs(ag) + H2O() 2. The KM for the decomposition of solid ammonium hydrogen sulfide to ammonia gas and hydrogen sulfide gas is 1.8 × 10-4 at 25°C. If ammonium hydrogen sulfide decomposes in a flask that already contains 0.020 M NHs(g), what are the equilibrium pressures of the products? 1?2?
Answer #1

Q1

K = 0.11

K = [acetate] / [ethanol][acid]

initially

[ethanol] = 0.81

[acid] = 0.81

[acetate] = 0

in equuilibirum

[ethanol] = 0.81 -x

[acid] = 0.81 - x

[acetate] = 0 +x

substitute in equilbirum

K = [acetate] / [ethanol][acid]

0.11 = x / (0.81-x)^2

0.11*(0.81^2 -2*0.81x +x^2) = x

9.09x = (0.81^2 -2*0.81x +x^2)

x^2 + (-2*0.81 - 9.09)x +(0.81^2) = 0

x = 0.0616

[ethyl acetate] = 0 +x = 0.0616 M

Q = (0.0616)/(0.81-0.0616)^2; which is 0.11, exactly K